OpenStudy (anonymous):

How do I get the inverse laplace transform of the following? I am completely lost... 6/(s+1)^3

OpenStudy (anonymous):

I'm tagging you again :-P @zepdrix

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

hmm, if i could recall the tables, it would help

OpenStudy (anonymous):

I have a table of transforms, but it does not have any helpful info. Well, not to me, anyway

OpenStudy (amistre64):

\[\frac{t^k}{k!}\to\frac{1}{s^{k+1}}\]

OpenStudy (amistre64):

in this case, you might wasnt to say u = (s+1) i dont think a partial fraction decomp would be useful ... but its been awhile

OpenStudy (anonymous):

Decomp didn't help.. I was stuck with a ()^2 and ()^3 again. Brb, going to try substitution

OpenStudy (amistre64):

recall the 6 = 3!

OpenStudy (amistre64):

#4 on the table seems to address it nicely

hartnn (hartnn):

no partial, that table doesn't have the required theorem

OpenStudy (amistre64):

hard to keep a focus when you have to flip between screens ....

OpenStudy (anonymous):

Okay I tried, but then we would totally ignore whatever is in the (), since the inverste transforms has no s's in. Or where are we going to substitute back again?

OpenStudy (amistre64):

\[\frac{6}{(s+1)^3}=3\frac{2!}{(s+1)^3}\]

OpenStudy (anonymous):

should I then just write (t+1) instead of t?

OpenStudy (amistre64):

consider: the laplace of say i believe so

OpenStudy (anonymous):

So, 3(t+1)^3?

OpenStudy (amistre64):

thats looking better yes; try ^2 tho

hartnn (hartnn):

use this : \(\huge L^{-1}F(s-a)=e^{at}f(t)\)

OpenStudy (amistre64):

yeah, theres in e in there

hartnn (hartnn):

so, from your Q, \(\huge e^{-t}[L^{-1}\dfrac{6}{s^3}]\) now use the table.

OpenStudy (anonymous):

So it's\[e^{-t}3t^2\]?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

Thanks! I'll do some more problems like these, I'm really struggling

hartnn (hartnn):

practice makes the man perfect! :)

OpenStudy (anonymous):

Thanks a lot Goku!

hartnn (hartnn):

welcome ^_^